# MemAscend: System Memory Optimization for SSD-Offloaded LLM Fine-Tuning

### v2

@@ I Introduction
Broadly, LLM training consumes two memory classes: (1) static memory, which includes model weights and optimizer states, and (2) residual states, such as activations and intermediate values [8]. For residual states, activations dominate due to their linear growth with batch size and sequence length [9], and several techniques have been proposed to mitigate the above memory pressure. For instance, to lower activation memory, gradient checkpointing [10] is designed not to store every activation during the forward pass, but instead stores only a few checkpointed activations and recomputes those missed ones during backpropagation. The intermediate values, which are temporary results that persist across multiple operations, such as across multiple CUDA kernels [^1] CUDA (Compute Unified Device Architecture) kernels are functions, also basic execution unit, written in CUDA C/C++ that run on the GPU. [11], significantly contribute to peak memory usage. To lower the size of intermediate values, methods like Liger-Kernel [12] fuse operations (e.g., RMSNorm, SwiGLU, Cross-Entropy [13, 14, 15]) into single Triton [16] kernels, eliminating transient buffers and launch overhead. Additionally, memory-efficient kernels like Flash-Attention [17] tile the computation to avoid materializing the entire attention matrix at once, thereby reducing the size of intermediate values. Other approaches, such as mixed-precision training [18] and related quantization techniques, can reduce the memory footprint of both activations and intermediate values by casting tensors to fp16 or bf16.
On the other hand, static memory is substantial because modern transformers contain billions of parameters, and optimizers such as Adam store both momentum and variance for each weight [19]. To lower the amount of static memory, quantization has been widely explored. For instance, 8-bit optimizers [20] and 4-bit optimizers [21] have shown promising results in reducing static memory consumption with minimal impact on convergence speed. On the other hand, Parameter-Efficient Fine-Tuning (PEFT) [22], especially Low-Rank Adaptation (LoRA) [23], significantly reduce the number of trainable parameters by focusing on low-rank matrices while keeping the pre-trained model frozen. Although methods like LoRA offer substantial computational and memory efficiency, they often require meticulous hyperparameter tuning, such as determining the rank and selecting which weights to train, to achieve optimal performance [24, 25]. Consequently, full-parameter training remains crucial for maximizing task-specific performance, particularly in scenarios where pre-trained models have not been exposed to domain-specific or private data. In such cases, the discrepancies between the target and pre-trained domains can be substantial, necessitating full-parameter training to bridge the gap effectively.
@@ II-A SSD offloading framework
To address GPU memory limitations in large-scale model training, heterogeneous memory systems that incorporate SSD offloading have emerged as a promising solution. These systems strategically relocate less frequently accessed data, primarily static model states, from expensive GPU memory to more abundant and cost-effective system memory or NVMe SSDs. ZeRO-Infinity [26] is a state-of-the-art (SOTA) SSD offloading system integrated into the DeepSpeed library [31]. Building on the ZeRO distributed training strategy [8, 32], which partitions model states across data-parallel processes to eliminate memory redundancy, ZeRO-Infinity extends the design by introducing an SSD offloading scheduler. This scheduler enables partitioned static model states to be offloaded first to CPU memory and subsequently to NVMe SSDs, as illustrated in Figure 1.
@@ III-C Peak Memory Bottlenecks in System Memory
To detect overflows during the overflow check, each value must be examined to determine whether it meets any of the following three conditions: infinity, negative infinity, or NaN. In PyTorch, this involves calling isinf() followed by any(). Internally, isinf() triggers isabs() to duplicate the tensor before performing comparisons, which creates a temporary copy and a Boolean tensor, resulting in a memory spike to about 2.25× the original size, as shown in Steps 2 of Figure 3. At Step 3, isnan() avoids isabs() but still produces a Boolean tensor, adding another 1.25× memory peak. For example, when training an 8-billion-parameter model, the flat buffer on the CPU consumes 29.91 GiB, but the overflow check pushes usage to 67.30 GiB. On the other hand, the sequence of operations in the overflow check introduces significant latency: (1) isabs(), (2) isinf(), (3) any(), (4) isnan(), and (5) any(), as shown in Steps 2 to 6 in Figure 3. This chain of steps adds roughly 5507 ms per iteration for an 8-billion-parameter model on the Configuration 1 machine described in Table II, accounting for 13.36% of the single iteration latency when training with a context length of 4096 and a batch size of 8, thereby substantially slowing down the training process. Since this check runs every iteration, the cumulative impact on performance is considerable. Thus, both the memory spike and the extra computation make gradient overflow checking a major bottleneck for large-scale models.
@@ III-D Filesystem Overhead in SSD offloading
TABLE I: Peak system memory usage comparison on different training model size across different training system approaches.
@@
III-E Motivation
Even though the SSD offloading technique offers a promising solution for training larger models on systems with limited resources, those observed system-memory bottlenecks intrinsic to SSD offloading lower scalability and erode cost-effectiveness. To understand the limit of existing SSD offloading techniques, a motivational experiment is conducted based on a machine with 24 GiB of GPU memory and a 128 GiB system memory limit, and results are summarized in Table I. Accordingly, the All-in-GPU approach, which stores everything on the GPU, supports training models with up to 1 billion parameters. By using ZeRO-Offload, which shifts data to CPU memory, the same system can handle models with up to 3 billion parameters. With the SOTA SSD offloading strategy, ZeRO-Infinity, it can train models with up to 8 billion parameters. It can be observed that, even though the model size is increased, the system memory usage grows more drastically and soon reaches the system limits of commodity CPUs. According to the above observations, inefficiencies such as buffer pool fragmentation, unnecessary alignment for pinned memory allocations, and memory spikes during gradient overflow checks lead to significant waste of system memory and prevent the SOTA SSD offloading strategy from accommodating larger models.
To quantify the actual required and wasted system memory usage, another motivational experiments are conducted on Configuration 2, as mentioned in Table II. As shown in Figure 4, the system wastes an average of 55.7% of its memory across different models, making system memory the next bottleneck. These limitations reduce the scalability and cost-effectiveness of SSD offloading. This study directly addresses those inefficiencies. Reclaiming wasted system memory can either increase the maximum trainable model size on current hardware or reduce the memory required for existing workloads, improving economic efficiency. The recovered capacity can also be used to store additional offloaded activation checkpoints, supporting longer context windows or larger batch sizes and boosting throughput by increasing useful computation per data movement. Finally, minimizing secondary overheads such as I/O latency and overflow check delays further enhances overall efficiency and scalability.
@@ IV-B Adaptive Buffer Pool
To accommodate the above buffer requirements at the system memory level, a buffer pool design is included in the original SSD offloading architecture. Nonetheless, as shown in Figure 6a, since the original design fixed the buffer block size to the largest buffer component, the issue of international fragmentation emerges. In this study, as illustrated in Figure 6b, the adaptive buffer pool design is included to assign separate subpools to tensors of different shapes, avoiding the internal fragmentation that plagues the original monolithic pool in SSD offloading. According to the flow of fine-tuning, four pools are sufficient: one for embedding-related weights, one for feed-forward projections, one for the K and V projections (identical in grouped-query attention), and one for the Q and O projections. Embedding and LM-head weights share the same dimensions—vocabulary × hidden—while feed-forward projections keep the intermediate × hidden shape, so fixed pool dimensions eliminate wasted space. The resulting subgroup counts are $2$, $3N$, $2N$, and $2N$, respectively, which match the exact number of tensors and effectively remove internal fragmentation, with N denoting the number of transformer blocks that must be in flight due to prefetching. The scheme generalizes to multi-GPU deployments that employ parameter partitioning: each process loads only its shard, so per-process buffers shrink proportionally with the number of partitions, even though the overall system still reserves the same total capacity as a single-GPU setup. Thus, the adaptive buffer pool retains its efficiency across both single-card and distributed configurations.
(a) Monolithic buffer pool in ZeRO-Infinity
(b) Adaptive Buffer Pool in MemAscend
@@ IV-C Alignment Free Pinned Memory Allocation
As noted in Section III-B, PyTorch’s standard caching allocator applies a power-of-two alignment policy for pinned memory allocations, which leads to significant memory waste, particularly when dealing with large static buffers used in SSD offloading. To mitigate this, a custom C++ extension has been included in the proposed Memasend. It allocates CPU memory directly using posix_memalign, aligning only to 4096 bytes to meet DMA requirements. After allocation, the memory is page-locked and registered with all CUDA contexts using the CUDA API cudaHostRegister, with the cudaHostRegisterPortable flag. A CPU-pinned tensor is then created via the torch::from_blob API, enabling efficient GPU DMA transfers. This direct allocation and registration approach ensures that pinned memory buffers occupy only the required space plus minimal alignment, avoiding the internal fragmentation caused by power-of-two rounding. As a result, system memory usage drops significantly for large allocations such as the gradient partition flat buffer and the adaptive buffer pool.
@@ IV-D Fused Overflow Check Mechanism
According to the discussion in Section III-C, the SSD offloading system detects gradient overflows by invoking operations, such as isnan() and isinf(), and leads to peak memory usage when transitioning between these functions. Combining these checks into a single pass removes the intermediate tensors each call would otherwise create, lowering peak memory demand. Gradients are produced in fp16 during back-propagation, yet the partition’s flat buffer is retained in fp32 for accumulation, so fp16-level overflows must still be vetted against fp32 representations. Algorithm 1 illustrates the fused procedure: under IEEE 754, overflow manifests when all exponent bits equal 1; NaN values (Not a Number) add a non-zero mantissa, whereas positive or negative infinity has a zero mantissa. Line 2 defines this bit pattern, and lines 5–8 apply bitwise logic to flag any value with an all-ones exponent as overflowed. Each element is independent, enabling wide parallel execution; practical tests show roughly 97% efficiency for models spanning 1 B–32 B parameters. By preventing intermediate allocations and exploiting bit-level checks, the fused mechanism removes the overflow-detection bottleneck at scale.
@@
[…]
@@
Figure 8: Peak system memory usage comparison for training Qwen2.5-7B between ZeRO-Infinity and MemAscend.
@@ IV-E Direct NVMe Engine
The primary goal of the direct NVMe engine is to bypass the Linux filesystem overhead by using raw I/O to directly send asynchronous I/O (AIO) requests to the device driver via iouring or libaio. The overview of architecture and its workflow is shown in Figure 7. First, the SSD offloading system issues a tensor AIO request. The request context contains a unique tensor name (as a key), a data pointer, and an offset aligned to the tensor size. Next, the direct NVMe engine distributes this request among multiple thread workers, dividing the data into equal portions for balanced performance. Each thread consults the tensor location dictionary to determine which SSD device and logical block address (LBA) to use. If the tensor location dictionary has not recorded the tensor yet, it consults the location allocator for SSD space. This step ensures that the data are horizontally partitioned across SSDs, effectively implementing striping in place of Software RAID 0. The shared device information structure, implemented via shared memory, guarantees that multiple processes do not allocate overlapping SSD regions. After each thread obtains its data pointer, offset, LBA, and raw SSD device path, it sends a raw AIO request through libaio. By applying load balancing, stripping, and multi-threading, the Direct NVMe Engine reduces latency and achieves higher throughput compared to approaches that rely on filesystem-based AIO.

### v3

@@ I Introduction
Broadly, LLM training consumes two memory classes: (1) static memory, which includes model weights and optimizer states, and (2) residual states, such as activations and intermediate values [8]. For residual states, activations dominate due to their linear growth with batch size and sequence length [9], and several techniques have been proposed to mitigate the above memory pressure. For instance, to lower activation memory, gradient checkpointing [10] is designed not to store every activation during the forward pass, but instead stores only a few checkpointed activations and recomputes those missed ones during backpropagation. The intermediate values, which are temporary results that persist across multiple operations, such as across multiple CUDA kernels [^1] CUDA (Compute Unified Device Architecture) kernels are functions, also basic execution units, written in CUDA C/C++ that run on the GPU. [11], significantly contribute to peak memory usage. To lower the size of intermediate values, methods like Liger-Kernel [12] fuse operations (e.g., RMSNorm, SwiGLU, Cross-Entropy [13, 14, 15]) into single Triton [16] kernels, eliminating transient buffers and launch overhead. Additionally, memory-efficient kernels like Flash-Attention [17] tile the computation to avoid materializing the entire attention matrix at once, thereby reducing the size of intermediate values. Other approaches, such as mixed-precision training [18] and related quantization techniques, can reduce the memory footprint of both activations and intermediate values by casting tensors to fp16 or bf16.
On the other hand, static memory is substantial because modern transformers contain billions of parameters, and optimizers such as Adam store both momentum and variance for each weight [19]. To lower the amount of static memory, quantization has been widely explored. For instance, 8-bit optimizers [20] and 4-bit optimizers [21] have shown promising results in reducing static memory consumption with minimal impact on convergence speed. On the other hand, Parameter-Efficient Fine-Tuning (PEFT) [22], especially Low-Rank Adaptation (LoRA) [23], significantly reduces the number of trainable parameters by focusing on low-rank matrices while keeping the pre-trained model frozen. Although methods like LoRA offer substantial computational and memory efficiency, they often require meticulous hyperparameter tuning, such as determining the rank and selecting which weights to train, to achieve optimal performance [24, 25]. Consequently, full-parameter training remains crucial for maximizing task-specific performance, particularly in scenarios where pre-trained models have not been exposed to domain-specific or private data. In such cases, the discrepancies between the target and pre-trained domains can be substantial, necessitating full-parameter training to bridge the gap effectively.
@@ II-A SSD offloading framework
To address GPU memory limitations in large-scale model training, heterogeneous memory systems that incorporate SSD offloading have emerged as a promising solution. These systems strategically relocate static model states, where tensor sizes vary dynamically due to mixed-precision storage formats, layer heterogeneity, and activation materialization boundaries. These tensors are not always needed during different stages of computation within a single training step, allowing them to be moved from expensive GPU memory to more abundant and cost-effective system memory or NVMe SSDs when not needed. The SSD acts as expanded but slower memory for storing data. Unlike traditional storage systems, where hot and cold data exhibit clear access-frequency disparities, SSD offloading during training accesses all offloaded tensors once per iteration. Therefore, all such data is effectively hot at every step. ZeRO-Infinity [26] is a state-of-the-art (SOTA) SSD offloading system integrated into the DeepSpeed library [31]. Building on the ZeRO distributed training strategy [8, 32], which partitions model states across data-parallel processes to eliminate memory redundancy, ZeRO-Infinity extends the design by introducing an SSD offloading scheduler. This scheduler enables partitioned static model states to be offloaded first to CPU memory and subsequently to NVMe SSDs, as illustrated in Figure 1.
@@ II-B3 LoHan
To clarify the positioning of MemAscend relative to contemporary SSD offloading techniques, Table I provides a comparative summary.
TABLE I: Comparison of MemAscend with recent SSD offloading techniques.
Technique Primary Scope and Memory Efficiency Hardware
ZeRO-Infinity General distributed offloading, but it suffers from memory fragmentation and overhead spikes. Commodity (Multi-GPU)
LoHan Minimizes host memory usage through integrated SSD activation offloading and employs holistic scheduling for optimal GPU utilization in single-GPU scenarios. Commodity (Single-GPU)
Smart-Infinity Near-storage processing for optimizer updates, but it does not address host DRAM bottlenecks. NSP-enabled SSD
MemAscend Optimizes general distributed fine-tuning and resolves ZeRO-Infinity’s inefficiencies. Commodity (Multi-GPU)
@@ III-C Peak Memory Bottlenecks in System Memory
To detect overflows during the overflow check, each value must be examined to determine whether it meets any of the following three conditions: infinity, negative infinity, or NaN. In PyTorch, this involves calling isinf() followed by any(). Internally, isinf() triggers isabs() to duplicate the tensor before performing comparisons, which creates a temporary copy and a Boolean tensor, resulting in a memory spike to about 2.25× the original size, as shown in Steps 2 of Figure 3. At Step 3, isnan() avoids isabs() but still produces a Boolean tensor, adding another 1.25× memory peak. For example, when training an 8-billion-parameter model, the flat buffer on the CPU consumes 29.91 GiB, but the overflow check pushes usage to 67.30 GiB. On the other hand, the sequence of operations in the overflow check introduces significant latency: (1) isabs(), (2) isinf(), (3) any(), (4) isnan(), and (5) any(), as shown in Steps 2 to 6 in Figure 3. This chain of steps adds roughly 5507 ms per iteration for an 8-billion-parameter model on the Configuration 1 machine described in Table III, accounting for 13.36% of the single iteration latency when training with a context length of 4096 and a batch size of 8, thereby substantially slowing down the training process. Since this check runs every iteration, the cumulative impact on performance is considerable. Thus, both the memory spike and the extra computation make gradient overflow checking a major bottleneck for large-scale models.
@@ III-D Filesystem Overhead in SSD offloading
III-E Motivation
Even though the SSD offloading technique offers a promising solution for training larger models on systems with limited resources, those observed system-memory bottlenecks intrinsic to SSD offloading lower scalability and erode cost-effectiveness. To understand the limit of existing SSD offloading techniques, a motivational experiment is conducted based on a machine with 24 GiB of GPU memory and a 128 GiB system memory limit, and results are summarized in Table II. Accordingly, the All-in-GPU approach, which stores everything on the GPU, supports training models with up to 1 billion parameters. By using ZeRO-Offload, which shifts data to CPU memory, the same system can handle models with up to 3 billion parameters. With the SOTA SSD offloading strategy, ZeRO-Infinity, it can train models with up to 8 billion parameters. It can be observed that, even though the model size is increased, the system memory usage grows more drastically and soon reaches the system limits of commodity CPUs. According to the above observations, inefficiencies such as buffer pool fragmentation, unnecessary alignment for pinned memory allocations, and memory spikes during gradient overflow checks lead to significant waste of system memory and prevent the SOTA SSD offloading strategy from accommodating larger models.
To quantify the actual required and wasted system memory usage, another motivational experiments are conducted on Configuration 2, as mentioned in Table III. As shown in Figure 4, the system wastes an average of 55.7% of its memory across different models, making system memory the next bottleneck. These limitations reduce the scalability and cost-effectiveness of SSD offloading. This study directly addresses those inefficiencies. Reclaiming wasted system memory can either increase the maximum trainable model size on current hardware or reduce the memory required for existing workloads, improving economic efficiency. The recovered capacity can also be used to store additional offloaded activation checkpoints, supporting longer context windows or larger batch sizes and boosting throughput by increasing useful computation per data movement. Finally, minimizing secondary overheads such as I/O latency and overflow check delays further enhances overall efficiency and scalability.
TABLE II: Peak system memory usage comparison on different training model size across different training system approaches.
@@ III-E Motivation
(a) Monolithic buffer pool in ZeRO-Infinity
(b) Adaptive Buffer Pool in MemAscend
@@ IV-B Adaptive Buffer Pool
To accommodate the above buffer requirements at the system memory level, a buffer pool design is included in the original SSD offloading architecture. Nonetheless, as shown in Figure 6(a), since the original design fixed the buffer block size to the largest buffer component, the issue of international fragmentation emerges. In this study, as illustrated in Figure 6(b), the adaptive buffer pool design is included to assign separate subpools to tensors of different shapes, avoiding the internal fragmentation that plagues the original monolithic pool in SSD offloading. According to the flow of fine-tuning, four pools are sufficient: one for embedding-related weights, one for feed-forward projections, one for the K and V projections (identical in grouped-query attention), and one for the Q and O projections. Embedding and LM-head weights share the same dimensions—vocabulary × hidden—while feed-forward projections keep the intermediate × hidden shape, so fixed pool dimensions eliminate wasted space. The resulting subgroup counts are $2$, $3N$, $2N$, and $2N$, respectively, which match the exact number of tensors and effectively remove internal fragmentation, with N denoting the number of transformer blocks that must be in flight due to prefetching. In this study, the adaptive buffer pool operates entirely on top of PyTorch’s allocator, modifying only logical buffer management. It preserves PyTorch’s safety and compatibility while reducing pinned-memory pressure, scaling effectively to multi-GPU training.
Notably, the management cost of the adaptive buffer pool is kept minimal by following the same underlying allocation approach as ZeRO-Infinity. This approach first allocates a single large, sequential monolithic tensor and then maintains metadata to manage which portion is assigned to each buffer. This design minimizes the allocation overhead that might occur with multi-pool management. When a buffer request occurs, such as when a tensor needs to be fetched from the SSD and copied to the CPU, a hashtable records a unique identification key. This key maps to the data structure containing the buffer’s metadata, which includes its size and the index of its portion within the large monolithic tensor. Since the number of distinct buffers is typically small (fewer than 100 in most mainstream open-source models, depending on the configuration), the complexity overhead of managing this hashtable metadata is negligible in terms of both computation and memory usage. Furthermore, this scheme efficiently generalizes to multi-GPU deployments that use parameter partitioning. In such a setup, each process loads only its own parameter shard, causing the per-process buffers to shrink proportionally with the number of partitions. Thus, the adaptive buffer pool maintains its efficiency in both single-GPU and distributed configurations.
@@ IV-C Alignment Free Pinned Memory Allocation
As noted in Section III-B, PyTorch’s standard caching allocator applies a power-of-two alignment policy. This policy introduces substantial memory waste when managing the large and static tensors essential for SSD offloading. These tensors are allocated once at initialization and persist until training completes, and they are major contributors to overhead when the power-of-two alignment is enforced due to their large size. These critical, parameter-related buffers include: (1) the gradient flat buffer for holding GPU-computed gradients during the backward pass, (2) all buffers within the parameter buffer pool used for prefetching model parameters from SSD to GPU, and (3) the optimizer state buffer that facilitates fetching and updating optimizer state subgroups on the CPU. To mitigate memory waste in these buffers, MemAscend incorporates a custom C++ extension to bypass this behavior and ensure memory efficiency through a precise, three-step process.
Direct Memory Allocation. First, the extension directly allocates CPU memory using posix_memalign. Unlike PyTorch’s default allocator, it aligns memory only to a 4096-byte boundary, which is the minimum required for Direct Memory Access (DMA) operations, thus eliminating the overhead from power-of-two rounding. Page-Locking and CUDA Registration. Next, the allocated memory is page-locked and registered with all available CUDA contexts via the cudaHostRegister API, using the cudaHostRegisterPortable flag. This step makes the memory block accessible to all GPUs for high-speed transfers. Pinned Tensor Creation. Finally, a CPU-pinned tensor is created from this memory block using the torch::from_blob API. This allows the PyTorch engine to use the buffer for efficient GPU DMA transfers without additional memory copies or the need to involve the CPU.
The aforementioned direct allocation and registration approach ensures that critical pinned buffers, such as the gradient partition flat buffer and the parameter buffer pool, occupy only the required space plus minimal alignment. Notably, this custom method is applied to static, large offloading buffers. Arbitrary, short-lived, or frequently allocated tensors still rely on PyTorch’s standard caching allocator to preserve the original memory management design choice. This strategy allows MemAscend to significantly reduce static overhead without introducing fragmentation risks that could compromise the system’s dynamic scalability. Furthermore, since changing the allocation size does not add any extra operations to the memory management workflow, the alignment-free pinned memory allocation introduces no additional overhead compared to the standard caching allocator.
@@ IV-D Fused Overflow Check Mechanism
According to the discussion in Section III-C, the SSD offloading system detects gradient overflows by invoking operations, such as isnan() and isinf(), and leads to peak memory usage when transitioning between these functions. Combining these checks into a single pass removes the intermediate tensors each call would otherwise create, lowering peak memory demand. Gradients are produced in fp16 during back-propagation, yet the partition’s flat buffer is retained in fp32 for accumulation, so fp16-level overflows must still be vetted against fp32 representations. Algorithm 1 illustrates the fused procedure: under IEEE 754, overflow manifests when all exponent bits equal 1; NaN values (Not a Number) add a non-zero mantissa, whereas positive or negative infinity has a zero mantissa. Line 2 defines this bit pattern, and lines 5–8 apply bitwise logic to flag any value with an all-ones exponent as overflowed. Each element is independent, enabling wide parallel execution; practical tests show roughly 97% efficiency for models spanning 1 B–32 B parameters. By preventing intermediate allocations and exploiting bit-level checks, the fused mechanism removes the overflow-detection bottleneck at scale. The proposed fused overflow check mechanism also generalizes efficiently to multi-GPU deployments, where each GPU process manages its own partition and performs the fused check without synchronization. The fused overflow check uses an OpenMP (Open Multi-Processing)-based parallel implementation, which introduces minor runtime overhead due to thread management, synchronization, and scheduling. However, these costs are negligible compared to the overall performance improvement, as the acceleration achieved effectively amortizes the OpenMP overhead.
@@
7:  if $(bits\,\&\,\texttt{EXP\_MASK})=\texttt{EXP\_MASK}$ then
@@
Figure 8: Peak system memory usage comparison for training Qwen2.5-7B between ZeRO-Infinity, MemAscend and theoretical minimal requirements.
@@ IV-E Direct NVMe Engine
The primary goal of the direct NVMe engine is to bypass the Linux filesystem overhead by using raw I/O to directly send asynchronous I/O (AIO) requests to the device driver via iouring or libaio. The overview of architecture and its workflow is shown in Figure 7. First, the SSD offloading system issues a tensor AIO request. The request context contains a unique tensor name (as a key), a data pointer, and an offset aligned to the tensor size. Next, the direct NVMe engine distributes this request among multiple thread workers, dividing the data into equal portions for balanced performance. Each thread consults the tensor location dictionary to determine which SSD device and logical block address (LBA) to use. If the tensor location dictionary has not recorded the tensor yet, it consults the location allocator for SSD space. This step ensures that the data are horizontally partitioned across SSDs, effectively implementing striping in place of Software RAID 0. The shared device information structure, implemented via shared memory, guarantees that multiple processes do not allocate overlapping SSD regions. Although this requires the overhead of accessing a shared offset counter for the initial allocation of a new, unrecorded tensor, the cost is negligible because it is a simple shared memory integer operation that occurs only once per tensor. After each thread obtains its data pointer, offset, LBA, and raw SSD device path, it sends a raw AIO request through libaio. By applying load balancing, stripping, and multi-threading, the Direct NVMe Engine reduces latency and achieves higher throughput compared to approaches that rely on filesystem-based AIO.